# The orbital evolution of the tidally stripped star and disk-driven stable mass transfer for QPEs in GSN069

### 13 Feb 2024 v2

@@
Abstract
@@ 2.1 Mass transfer
$R=9\times 10^{8}\left[1-\left(\frac{M}{M_{\mathrm{ch}}}\right)^{4/3}\right]^{1/2}\left(\frac{M}{M_{\odot}}\right)^{-1/3}\mathrm{~{}cm},$ (4)
@@ 2.3 Orbital Evolution for the WD
Here we show an example of the evolution under mass transfer and gravitational radiation in Fig. 1. We adopt possible parameters of GSN 069: $P_{b}=9\mathrm{~{}hr}$, $M_{h}=5\times 10^{5}M_{\odot}$ and $M=0.21M_{\odot}$. Then $e=0.968$ can be obtained from Eq. (4) with $\beta\approx\beta_{0}$. The initial $\beta$ is 0.50001. The orbital parameters and the mass and radius of the WD are updated each time the WD passes through the periastron. We use $\langle\dot{r_{p}}\rangle$ and $\langle\dot{e}\rangle$ to evolve orbit, because $\langle\dot{a}/a\rangle$ is too high to properly represent orbital changes in the above method when $1-e$ is very small. Evolution will stop when $e$ is greater than 1.
@@ 2.4 Orbital Evolution for the main-sequence star
We show an example in Fig. 3. The initial orbital parameters are $M_{h}=5\times 10^{5}M_{\odot}$, $M=0.1M_{\odot}$, $\beta=0.50004$ and $P_{b}=40\mathrm{~{}hr}$. The radius is obtained by $R=1.06(M/M_{\odot})^{0.945}R_{\odot}$ (Demircan & Kahraman, 1991). Then the eccentricity can be obtained $e=0.912$.
@@ 3.1 Drag force by the accretion disk
$F(\mathcal{M})=\begin{cases}\frac{1}{2}\ln\left(\frac{1+\mathcal{M}}{1-\mathcal{M}}\right)-\mathcal{M},&0<\mathcal{M}<1,\\ \frac{1}{2}\ln\left(1-\frac{1}{\mathcal{M}^{2}}\right)+\ln\left(\frac{r}{R}\right),&\mathcal{M}>1.\end{cases}$ (18)
@@ 3.2.1 Stable mass transfer
Figure 5: Timescales with true anomaly for the hydrodynamic force and dynamical friction for a WD with $M=0.21M_{\odot}$, $P_{b}=9\mathrm{~{}hr}$ and $e=0.95$. The upper panel is for the prograde star and the lower panel is for the retrograde star. Solid lines and dashed lines represent the timescales of the hydrodynamic force and dynamical friction, respectively. Different colors represent the different accretion rate of the disk.
We consider an $\alpha$ standard disk with $\alpha=0.1$ and $M_{h}=5\times 10^{5}\mathrm{~{}M_{\odot}}$. The density and sound speed can be calculated for a given accretion rate $\dot{M}_{Acc}$ in the unit of Eddington accretion rate $\dot{M}_{Edd}$ (radiation efficiency $\eta=0.1$). As shown in Fig. 5, the hydrodynamic force is always stronger than the dynamical friction in all orbital phases, thus dominant on the effect by disk. We only consider the hydrodynamic force in the following calculation.
Figure 6: Secular change of orbital parameters caused by hydrodynamic force for a prograde WD with $M=0.21M_{\odot}$ and $P_{b}=9\mathrm{~{}hr}$ under different accretion rate. The top panel shows the change of $\beta$, specific angular momentum, and specific orbital energy caused by drag force, it also shows the effect on $\beta$ caused by gravitational radiation for comparison. The dashed black line represents the point with $\dot{M}_{Acc}$ of GSN 069 and $e=0.95$, it predicts $\dot{\beta}/\beta\approx-3.5\times 10^{-4}\mathrm{~{}yr}^{-1}$(see texts). The bottom panel represents the change of the $a$ and $e$ relative to $\beta$.
@@
Here we adopt the results in Miniutti et al. (2023b). Assuming radiation efficiency $\eta=0.1$, the accreted mass is about $3.7\times 10^{-8}\mathrm{~{}M_{\odot}}$, corresponding to $\dot{M}/M\approx-1.7\times 10^{-4}\mathrm{yr}^{-1}$ and $\beta\approx 0.5005$ with $M=0.21\mathrm{~{}M_{\odot}}$. The QPEs intensity is roughly halved per year, corresponding to $\dot{\Delta M}/\Delta M=-0.5\mathrm{~{}yr}^{-1}$ in Eq. (2), which requires $\dot{\beta}/\beta=-2\times 10^{-4}\mathrm{~{}yr}^{-1}$.
The change of $\beta$ induced by mass transfer is $\langle\dot{\beta}/\beta\rangle_{\dot{M}}\approx 1.1\times 10^{-4}\mathrm{~{}yr}^{-1}$ with $\zeta=-1/3$. The bolometric luminosity of TDE is about $10^{44}\mathrm{~{}erg/s}$ corresponding to $\sim 1.6\dot{M}_{Edd}$. Assuming $e=0.95$, it predicts $\langle\dot{\beta}/\beta\rangle_{disk}\approx-3.5\times 10^{-4}\mathrm{~{}yr}^{-1}$ as shown in Fig. 6. Both of them give $\dot{\beta}/\beta\approx-2.4\times 10^{-4}\mathrm{~{}yr}^{-1}$, which is consistent with the required value. Consequently, the hydrodynamic drag force can explain the decaying of the intensity of QPEs from GSN 069.
@@ 5 Summary
As the accretion rate of the fallback disk gradually decreases, the drag force by the disk will decrease until it is negligible. Then the mass transfer will start to be unstable eventually the WD will escape the SMBH. Continued observations of GSN 096 will help us to understand the interaction of the disk with the star inside it and to test our model. By accounting for the effect of the mass transfer, the evolutionary path of the stripped star is different from the evolutionary only with gravitational wave radiation. The criterion of the unstable mass transfer is changed and the fate of the stripped star is to escape the SMBH rather than be disrupted by it. In this case, however, mass transfer remains unstable.

### 05 Mar 2024 v3

@@
11institutetext: Department of Physics, Huazhong University of Science and Technology, Wuhan 430074, China
@@
The orbital evolution of the tidally stripped star and disk-driven stable mass transfer for QPEs in GSN 069
Di Wang
@@ 2.1 Mass transfer
$R=9\times 10^{8}\left[1-\left(\frac{M}{M_{\mathrm{ch}}}\right)^{4/3}\right]^{1/2}\left(\frac{M}{M_{\odot}}\right)^{-1/3}\mathrm{\leavevmode\nobreak\ cm},$ (4)
@@ 2.3 Orbital Evolution for the WD
Here we show an example of the evolution under mass transfer and gravitational radiation in Fig. 1. We adopt possible parameters of GSN 069: $P_{b}=9\mathrm{\leavevmode\nobreak\ hr}$, $M_{h}=5\times 10^{5}M_{\odot}$ and $M=0.21M_{\odot}$. Then $e=0.968$ can be obtained from Eq. (4) with $\beta\approx\beta_{0}$. The initial $\beta$ is 0.50001. The orbital parameters and the mass and radius of the WD are updated each time the WD passes through the periastron. We use $\langle\dot{r_{p}}\rangle$ and $\langle\dot{e}\rangle$ to evolve orbit, because $\langle\dot{a}/a\rangle$ is too high to properly represent orbital changes in the above method when $1-e$ is very small. Evolution will stop when $e$ is greater than 1.
@@ 2.4 Orbital Evolution for the main-sequence star
We show an example in Fig. 3. The initial orbital parameters are $M_{h}=5\times 10^{5}M_{\odot}$, $M=0.1M_{\odot}$, $\beta=0.50004$ and $P_{b}=40\mathrm{\leavevmode\nobreak\ hr}$. The radius is obtained by $R=1.06(M/M_{\odot})^{0.945}R_{\odot}$ (Demircan & Kahraman, 1991). Then the eccentricity can be obtained $e=0.912$.
@@ 3.1 Drag force by the accretion disk
[…]
@@ 3.2.1 Stable mass transfer
Figure 5: Timescales with true anomaly for the hydrodynamic force and dynamical friction for a WD with $M=0.21M_{\odot}$, $P_{b}=9\mathrm{\leavevmode\nobreak\ hr}$ and $e=0.95$. The upper panel is for the prograde star and the lower panel is for the retrograde star. Solid lines and dashed lines represent the timescales of the hydrodynamic force and dynamical friction, respectively. Different colors represent the different accretion rate of the disk.
We consider an $\alpha$ standard disk with $\alpha=0.1$ and $M_{h}=5\times 10^{5}\mathrm{\leavevmode\nobreak\ M_{\odot}}$. The density and sound speed can be calculated for a given accretion rate $\dot{M}_{Acc}$ in the unit of Eddington accretion rate $\dot{M}_{Edd}$ (radiation efficiency $\eta=0.1$). As shown in Fig. 5, the hydrodynamic force is always stronger than the dynamical friction in all orbital phases, thus dominant on the effect by disk. We only consider the hydrodynamic force in the following calculation.
Figure 6: Secular change of orbital parameters caused by hydrodynamic force for a prograde WD with $M=0.21M_{\odot}$ and $P_{b}=9\mathrm{\leavevmode\nobreak\ hr}$ under different accretion rate. The top panel shows the change of $\beta$, specific angular momentum, and specific orbital energy caused by drag force, it also shows the effect on $\beta$ caused by gravitational radiation for comparison. The dashed black line represents the point with $\dot{M}_{Acc}$ of GSN 069 and $e=0.95$, it predicts $\dot{\beta}/\beta\approx-3.5\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$(see texts). The bottom panel represents the change of the $a$ and $e$ relative to $\beta$.
@@
Here we adopt the results in Miniutti et al. (2023b). Assuming radiation efficiency $\eta=0.1$, the accreted mass is about $3.7\times 10^{-8}\mathrm{\leavevmode\nobreak\ M_{\odot}}$, corresponding to $\dot{M}/M\approx-1.7\times 10^{-4}\mathrm{yr}^{-1}$ and $\beta\approx 0.5005$ with $M=0.21\mathrm{\leavevmode\nobreak\ M_{\odot}}$. The QPEs intensity is roughly halved per year, corresponding to $\dot{\Delta M}/\Delta M=-0.5\mathrm{\leavevmode\nobreak\ yr}^{-1}$ in Eq. (2), which requires $\dot{\beta}/\beta=-2\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$.
The change of $\beta$ induced by mass transfer is $\langle\dot{\beta}/\beta\rangle_{\dot{M}}\approx 1.1\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$ with $\zeta=-1/3$. The bolometric luminosity of TDE is about $10^{44}\mathrm{\leavevmode\nobreak\ erg/s}$ corresponding to $\sim 1.6\dot{M}_{Edd}$. Assuming $e=0.95$, it predicts $\langle\dot{\beta}/\beta\rangle_{disk}\approx-3.5\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$ as shown in Fig. 6. Both of them give $\dot{\beta}/\beta\approx-2.4\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$, which is consistent with the required value. Consequently, the hydrodynamic drag force can explain the decaying of the intensity of QPEs from GSN 069.